# A Data Generation Perspective to the Mechanism of In-Context Learning

### v2

@@ 4.1 Bayesian Selection in the Function Learning Framework
The Bayesian perspective can be found in the function learning framework originally utilized for the skill learning mechanism. Typically, we illustrate the underlying Bayesian selection in the function learning framework, indicating the intrinsic connection between the two statistical frameworks. According to Ahuja et al. (2023), the transformers pre-trained on the data generated from diverse function classes exhibit improved function-fitting ability across all the pre-training function classes. To identify the best-fit solution among the whole function class, the function selection process implements a Bayesian optimal selection. More details can be found in Section 3.2. Notably, instead of the original Bayesian inference framework only selecting the pre-training data generation functions, the function selection scope is enlarged, including all the unseen functions from the same function class with the pre-training functions.
4.2 Extending the Bayesian Inference Framework for Skill Learning
@@
We then illustrate the possibility of extending the Bayesian inference framework to understand the skill learning mechanism to capture new data generation functions from the in-context data via relaxing the particular assumption. One important assumption in the Bayesian inference framework (Xie et al., 2021) is that all ICL demonstrations should be generated with the same latent concept. Nonetheless, this strong assumption may not be held in practice. For instance, one demonstration sample discusses the topic of sociology but another one is relevant to cardiology, the data generation function for these two domains should be rather different. Inspired by the high compositionality nature of language data, Hahn & Goyal (2023) came up with an information-theoretic bound showing that ICL performance can be improved given more unique compositional structures in pre-training data, therefore skill learning ability can appear by combining compositionality structures, in pre-training data, to infer the data generation function of ICL demonstrations.
@@
Empirical evidence shows that, given an input-label pair of two semantically unrelated concepts, e.g., mapping sports to animals, Rong (2021); Wei et al. (2023) still observe a satisfactory performance with the increasing model scale, indicating that the LLM can retrieve multiple concepts and combine them as a new data generation process. Feng & Steinhardt (2023) interpret the combination with a binding mechanism with an internal function vector to recognize the input feature and bind it to the corresponding label.
@@
Swaminathan et al. (2023) proposes another way to extend the existing Bayesian framework for skill learning via replacing the original HMM model into the clone-structured causal graph (CSCG) (George et al., 2021; Dedieu et al., 2019). The major difference is that the CSCG considers a learnable emission matrix, which determines the probability of observing a particular output given each hidden state in the model. A relevant transition matrix as the concept is retrieved, similar to the Bayesian inference (Xie et al., 2021). The hidden states for each token can then be obtained given the particular relevant template. The LLM then learns the suitable emission matrix, providing the best-fit mapping from the hidden states to the observed token.
5 Empirical Investigation On Skill Recognition and Skill Learning
@@
In this section, we exhibit more empirical analyses revolving around skill recognition and skill learning abilities. In contrast to the mechanism analysis that focuses on whether the ICL can learn new in-context data generation functions or not, empirical evidence in this section indicates that it is highly likely that LLMs exhibit both skill recognition and skill learning abilities of various levels, instead of an all-or-nothing conclusion. We first discuss how the LLM jointly obtains both abilities during the pre-training stage in Section 5.1. Specifically, the origin of both abilities is determined by the pre-training data distribution (Chan et al., 2022) and the model scale (Wei et al., 2023; Pan, 2023). We then investigate how the LLM effectively utilizes the obtained abilities during the ICL inference stage in Section 5.2. Typically, the LLM exhibits varying degrees of usage on those two abilities according to tasks with different difficulties. The unique strengths and weaknesses of each ability are shown in Section 5.3.
5.1 Origin of Skills
@@
Analyses are first conducted focusing on how those abilities are developed along the pre-training procedure. (Bietti et al., 2023) observe that the skill recognition ability is obtained early in the pre-training procedure, while the skill learning ability is developed much later. However, Singh et al. (2023) shows that the obtained skill learning ability gradually vanishes after over-training and is replaced by the skill recognition ability. Such observation indicates that skill learning is a transient ability that may disappear when the model is over-trained rather than a persistent one which can be kept once obtained. The reason can be attributed to the pre-training data distribution (Chan et al., 2022) where the task learning ability degrades if the pre-training data follows a uniform, i.i.d distribution. Nonetheless, such degradation may not happen when the pre-training data follows a properly skewed Zipfian distribution. Chan et al. (2022) further emphasizes that the skill learning ability emerges when the pre-training data meets the following properties: (1) Skewed rank-frequency distributions: Dynamic contextual meaning does not uniform across data, instead, only a few meanings dominate with the long tail of other infrequent meanings. (2) Burstiness: Dynamic contextual meaning is not uniform across time, but appears in clusters. The reason why ICL ability can be obtained on such data distribution remains unclear. A potential explanation could be that the pre-training weight can only obtain the head meaning frequently appears while the long tail knowledge can only be obtained via ICL.
@@
Analyses are then conducted with a focus on the impact of the model scale. Pan (2023) illustrates that the skill recognition ability can be found across LLMs with different scales. In contrast, LLMs obtain better skill learning ability along with an increasingly larger scale. Similar observations can be found in (Wei et al., 2023) that the LLM can learn the flipped input-label mapping and override pre-training knowledge when the model scale is sufficiently large. (Fu et al., 2023c) provides the potential explanation for the above phenomenon where the good skill recognition ability serves as a necessity for developing the skill learning ability.
5.2 Application of Skills
After the LLM obtained the skill learning and skill recognition abilities during pre-training, we then investigate how the model utilizes both abilities for achieving satisfactory downstream task performance during the ICL inference stage. Overall, the behavior of the LLM is more consistent with the skill recognition mechanism on difficult tasks while observations aligned with skill learning are more common to see on easy tasks.
5.3 Advantage and Disadvantage of Skills
@@
Considering the intricate interplay of both abilities on different tasks, we further illustrate the strengths and weaknesses inherent in each ability. Skill learning ability can obtain new knowledge from the in-context data, and even over-ride the pre-training knowledge. It provides an easy way to update the knowledge on the specific application without requiring computationally heavy fine-tuning. Such ability has been successfully utilized in different LLM applications, e.g. model editing with ICL (Zheng et al., 2023). Nonetheless, the skill learning ability may fail as it can be easily distracted by irrelevant context (Shi et al., 2023). The failure reason found in (Tang et al., 2023) is that the input-label mapping favors the shortcut as the model scale increases. Skill recognition ability is insensitive to the new in-context pattern leading to the failure on the specification-heavy task (Peng et al., 2023) while it exhibits robustness to the incorrectness of label-demonstrations and other in-context noise (Webson & Pavlick, 2021). Based on the above discussion, we suggest a careful evaluation of LLM about each ability and select a desired one for the downstream task.
6 Future Directions
How the skill learning ability emerges during pre-training. The emergence of the skill learning ability can be partially attributed to the skewed rank-frequency distribution of pre-training corpora (Chan et al., 2022), and Reddy (2023) highlights the role of the induction head (Olsson et al., 2022), a particular attention head which explicitly searches for a prior occurrence of the current token in-context and copying the suffix as prediction. Moreover, the function class-based analysis (Raventos et al., 2023) illustrates that the transition from skill recognition to skill learning only happens given diverse enough tasks in pre-training corpora. It is interesting to explore how these factors collaboratively influence the emergence of skill learning.
Why ICL can only learn data generation function appeared during the pre-training? In section 3, we provide a comprehensive discussion on what function can be learned in context. Observations indicate that ICL can only learn the function within the pre-training data generation function class. Nonetheless, the causality of the pre-training data generation function to ICL remains unclear. Garg et al. (2022) proposes the research question as Can we train a model to in-context learn a certain function class but overlooks the effect of the pre-training data generation function class. Once we have a certain clue about causality, we can leverage the skill-learning ability in a more controllable and safe manner.
Extending existing findings to other capabilities of LLMs. Despite the very original ICL being for classification tasks, more capabilities are observed, e.g., step-by-step reasoning ability (Wei et al., 2022b) for reasoning and self-correction (Ganguli et al., 2023). A critical question is how we can extend the understanding frameworks introduced in this paper, particularly the data generation perspective, to more complicated LLMs’ capabilities. Some pioneering research has been done; Prystawski & Goodman (2023) extends the Bayesian inference framework to understand the effectiveness of the CoT prompt. Kadavath et al. (2022) focuses on the self-evaluation prompt showing that LLMs can accurately examine the correctness of their statements. We believe the introduced data generation perspective and two main understanding frameworks on ICL serve us the milestone to explore more intrinsic capacities of LLMs.
7 Conclusion
In this study, we introduce a novel data generation perspective and use it to understand the underlying mechanism driving the current success of ICL. We primarily focus on understanding the LLM’s ability of skill learning and skill recognition, and investigating whether or not ICL inference is capable of learning new data generation functions in context. Given the increasing significance of LLMs, enhancing our understanding of these underlying processes is crucial. Our work makes a step forward in this direction, hopefully, it will stimulate further investigation of this important topic.

### v3

@@ A.1 Bayesian Selection in the Function Learning Framework
The Bayesian perspective can be found in the function learning framework originally utilized for the skill learning mechanism. Typically, we illustrate the underlying Bayesian selection in the function learning framework, indicating the intrinsic connection between the two statistical frameworks. According to Ahuja et al. (2023), the transformers pre-trained on the data generated from diverse function classes exhibit improved function-fitting ability across all the pre-training function classes. To identify the best-fit solution among the whole function class, the function selection process implements a Bayesian optimal selection. More details can be found in Section 5.2. Notably, instead of the original Bayesian inference framework only selecting pre-training data generation functions, the function selection scope is enlarged, including all the unseen functions from the same function class with the pre-training functions.
A.2 Extending the Bayesian Inference Framework for Skill Learning
@@
We then illustrate the possibility of extending the Bayesian inference framework to understand the skill learning mechanism to capture new data generation functions from the in-context data via relaxing the particular assumption. One important assumption in the Bayesian inference framework Xie et al. (2021) is that all ICL demonstrations should be generated with the same latent concept. Nonetheless, this strong assumption may not be held in practice. For instance, one demonstration sample discusses the topic of sociology but another one is relevant to cardiology, the data generation function for these two domains should be rather different. Inspired by the high compositionality nature of language data, Hahn and Goyal (2023) came up with an information-theoretic bound showing that ICL performance can be improved given more unique compositional structures in pre-training data, therefore skill learning ability can appear by combining compositionality structures, in pre-training data, to infer the data generation function of ICL demonstrations.
@@
Empirical evidence shows that, given an input-label pair of two semantically unrelated concepts, e.g., mapping sports to animals, Rong (2021); Wei et al. (2023) still observe a satisfactory performance with the increasing model scale, indicating that the LLM can retrieve multiple concepts and combine them as a new data generation process. Feng and Steinhardt (2023) interpret the combination with a binding mechanism with an internal function vector to recognize the input feature and bind it to the corresponding label.
@@
Swaminathan et al. (2023) proposes another way to extend the existing Bayesian framework for skill learning via replacing the original HMM model into the clone-structured causal graph (CSCG) George et al. (2021); Dedieu et al. (2019). The major difference is that the CSCG considers a learnable emission matrix, which determines the probability of observing a particular output given each hidden state in the model. A relevant transition matrix as the concept is retrieved, similar to the Bayesian inference Xie et al. (2021). The hidden states for each token can then be obtained given the particular relevant template. The LLM then learns the suitable emission matrix, providing the best-fit mapping from the hidden states to the observed token.
Appendix B Empirical Investigation On Skill Recognition and Skill Learning
@@
In this section, we exhibit more empirical analyses revolving around skill recognition and skill learning abilities. In contrast to the mechanism analysis that focuses on whether the ICL can learn new in-context data generation functions or not, empirical evidence in this section indicates that it is highly likely that LLMs exhibit both skill recognition and skill learning abilities of various levels, instead of an all-or-nothing conclusion. We first discuss how the LLM jointly obtains both abilities during the pre-training stage in Section B.1. Specifically, the origin of both abilities is determined by the pre-training data distribution Chan et al. (2022a) and the model scale Wei et al. (2023); Pan (2023). Typically, the LLM exhibits varying degrees of usage on those two abilities according to tasks with different difficulties.
B.1 Origin of Skills
@@
Analyses are first conducted focusing on how those abilities are developed along the pre-training procedure. Bietti et al. (2023) observe that the skill recognition ability is obtained early in the pre-training procedure, while the skill learning ability is developed much later. However, Singh et al. (2023) shows that the obtained skill learning ability gradually vanishes after over-training and is replaced by the skill recognition ability. Such observation indicates that skill learning is a transient ability that may disappear when the model is over-trained rather than a persistent one which can be kept once obtained. The reason can be attributed to the pre-training data distribution Chan et al. (2022a) where the task learning ability degrades if the pre-training data follows a uniform, i.i.d distribution. Nonetheless, such degradation may not happen when the pre-training data follows a properly skewed Zipfian distribution. Chan et al. (2022a) further emphasizes that the skill learning ability emerges when the pre-training data meets the following properties: (1) Skewed rank-frequency distributions: Dynamic contextual meaning does not uniform across data, instead, only a few meanings dominate with the long tail of other infrequent meanings. (2) Burstiness: Dynamic contextual meaning is not uniform across time, but appears in clusters. The reason why ICL ability can be obtained on such data distribution remains unclear. A potential explanation could be that the pre-training weight can only obtain the head meaning frequently appears while the long tail knowledge can only be obtained via ICL.
@@
Analyses are then conducted with a focus on the impact of the model scale. Pan (2023) illustrates that the skill recognition ability can be found across LLMs with different scales. In contrast, LLMs obtain better skill learning ability along with an increasingly larger scale. Similar observations can be found in Wei et al. (2023) that the LLM can learn the flipped input-label mapping and override pre-training knowledge when the model scale is sufficiently large. Fu et al. (2023b) provides the potential explanation where the good skill recognition ability serves as a necessity for developing the skill learning ability.
Appendix C Skill Composition
We primarily focus on the skill learning ability where the ICL can learn a new data generation function, and skill recognition ability where the ICL utilizes the data generation function from pre-training data. Instead of focusing on the single data generation function, combining multiple data generation functions together can lead to a complicated data generation function. We named such capability as skill composition capability, helping the LLM to achieve a complicated task by combining a sequence of simple and basic steps. Arora and Goyal (2023) theoretically indicates the effectiveness of skill composition where the complicated task can exhibit emergent performance gain when all the decomposed basic skills improve linearly.
The discussions on skill composition are organized as follows. In Section C.1, we investigate the effectiveness of skill composition ability. In Section C.2, we analyze when the skill composition capability can work. In Section C.3, we further illustrate more discussion and real-world applications on the skill composition ability. Notably, the skill composition ability is complicated without a general data generation function framework so far. The skill-composition ability often requires to be elicited by specific-designed ICL prompts, e.g., Chain-of-Thought prompting (CoT) Wei et al. (2022b), Tree-of-thought Yao et al. (2023), and Graph-of-Thought Besta et al. (2023), which generates multiple intermediate steps before the final answer. Most following literature conducts analysis on the CoT prompt.
C.1 Effectiveness of Skill Composition
In this section, we investigate the effectiveness of skill composition ability. Feng et al. (2023) indicates that if the skill decomposition is applied, the LLM can be more expressive to describe more complicated problems, e.g., mathematical and decision-making problems. Li et al. (2023b); Yang et al. (2023) further demonstrate the data efficiency where the skill composition facilitates can learn complicated functions with a reduced sample complexity. Prystawski and Goodman (2023) attributes the above expressiveness and efficiency with the local structures in the training data generation function. Such locality enables to accurate inference on each intermediate step supported by the similar pre-training data generation function. In contrast, direct inference as a whole instead of each local steps are likely to fail requiring since such complicated data generation function does not appear during the pre-training stage. In summary, the skill composition ability of LLMs enhances their expressiveness and data efficiency for modeling complicated data generation function, building on the basis of locality data generation function from the pre-training data.
C.2 When Skill Composition Works
We demonstrate the effectiveness of the composition in Section C.1, however, it remains unknown whether the decomposed intermediate steps are well-organized aligning with human cognition. To examine the correctness of the LLM decomposition, the literature focuses on formal deductive reasoning tasks like math reasoning Ahn et al. (2024). It enables to conducting systematic and controllable analysis on each reasoning step with the unique correct answer.
LLMs are able to conduct correct decomposition on particular tasks, aligning with the ideal human reasoning process. Zhou et al. (2023) finds a theoretical criterion to identify when the LLM can implement the ideal decomposition. Typically, when the task can be described by a short RASP program Weiss et al. (2021), a programming language designed for the computational model of a Transformer, the LLM can achieve the correct decomposition. Similarly, Yao et al. (2021) demonstrates that the transformer can process correct decomposition on particular formal languages with hierarchical structure, e.g., $\text{Dyck}_{k}$ Chomsky and Schützenberger (1959). With a suitable decomposition, LLMs can easily solve arbitrary complicated problems Jelassi et al. (2023); Li and McClelland (2023).
Beyond those identified tasks, it remains many tasks where LLMs cannot conduct an ideal decomposition. The key underlying reason McCoy et al. (2023) is the gap between human cognition and the next-token prediction pre-training task, requiring to tackle problems sequentially greedily. Instead of a proper decomposition, a greedy shortcut can be obtained from standard training, which skips the particular step instead of a formal decomposition. Theoretical evidence on the existence of shortcuts can be found in Liu et al. (2022) on the semi-automaton reasoning task. Saparov and He (2022) indicates that the shortcut can easily select the wrong step, leading to an incomplete planning and subsequently an incorrect answer, leading to failure on complicated tasks Dziri et al. (2023). Such inherent failure is unavoidable as the transformer always finds a shortcut solution Liu et al. (2022) while impossibile to find the exact implementation of the semi-automaton reasoning requiring recurrent models of computation with shallow and non-recurrent architecture. On the contrary, the shortcut also shows its benefits, converting the original complicated reasoning problem with multiple hops into a simpler one with less hops Wu et al. (2023b); Saparov and He (2022), alleviating the performance degradation along with the increased hop.
In summary, the shortcut solution of LLMs can be a double-side sword to solve a compositional problem. Nonetheless, it remains no existing study on how the LLM acquires the decomposition capability from pre-training data. Notably, we focus on whether the LLM composition aligns with the human decomposition while the manually-conducted deduction rules may not be optimal. The optimal decomposition remains unknown.
C.3 More Discussions
Despite the above comprehensive understanding, there are more empirical studies on the skill composition ability from various perspectives as follows. Madaan and Yazdanbakhsh (2022) divides the CoT prompt into three key components: symbols, patterns, and text with distinct roles as follows: (1) The exact type of symbols does not matter. (2) The patterns are the template serving as a trigger helping to locate the correct concept (3) Text contains commonsense knowledge and meaning, leading to the ultimate success. Similarly, Wang et al. (2022) divides the CoT prompt into two key components: bridging objects (the key and necessary objects) and language templates. Interestingly, neither of them matters. In contrast, the relevance to the query and correct reasoning ordering matters.
More recently, Xu et al. (2024) challenges the skill compositional capability of LLMs, pointing out the failure on the sequential reasoning tasks. On the contrary, LLMs can perform well on simple composite tasks that can be easily separated into sub-tasks based on the inputs solely. The skill composition ability remains mysterious, requiring further analyses.
Appendix D Transformer architecture simplification
To facilitate analysis, many studies introduce necessary simplifications to the standard Transformer architecture. While all empirical analyses use the standard Transformer setup, theoretical analyses adopt a modified version without layer normalization. More detailed theoretical simplification can be found as follows.
  • Do not consider model architecture Xie et al. (2021); Hahn and Goyal (2023)
  • A single-layer linear attention Von Oswald et al. (2023); Ahn et al. (2023a); Mahankali et al. (2023)
  • A single-layer relu attention Fu et al. (2023a)
  • A single-layer softmax attention Zhao (2023); Zhang et al. (2023c); Ren and Liu (2023); Li et al. (2023a)
  • An L-layer linear attention Ahn et al. (2023a)
  • A single-layer linear attention with FFN Von Oswald et al. (2023)
  • A full transformer Akyürek et al. (2022); Cheng et al. (2023); Bai et al. (2023); Guo et al. (2023)
Appendix E Discussions
E.1 The Emergence Phenomenon On the ICL Generalization
Chan et al. (2022b) proposes an interesting perspective to characterize how the ICL generalizes to the test data based on the in-context samples. Observations exhibit that the larger LLMs can achieve rule-based generalization similarly with the SVM. The rule-based generalization makes decisions using a minimal set that is central to the category definition, disregarding less essential data, Nonetheless, induction heads mechanism with prefix match and copy are more aligned with examplar-based generalization like KNN. The reason why LLM can achieve rule-based generalization still remains unclear.
E.2 Advantages And Disadvantages of Skill Learning And Skill Recognition
@@
Skill learning mechanism can obtain new knowledge from the in-context pattern, and even over-ride the pre-training knowledge. It provides an easy way to update the knowledge on the specific application without requiring computational-heavy fine-tuning. Such ability has been successfully utilized in different LLM applications, e.g. model editing with ICL Zheng et al. (2023). Nonetheless, the skill learning mechanism may fail as it can be easily distracted by irrelevant context Shi et al. (2023). The failure reason found in Tang et al. (2023) is that the input-label mapping is more to be the shortcut as the model scale increases. Skill recognition mechanism is insensitive to the new in-context pattern leading to the failure on the specification-heavy task Peng et al. (2023) while it exhibits robustness to the incorrectness of label-demonstrations and other in-context noise Webson and Pavlick (2021). For instance, the skill recognition mechanism can perform well in a noisy setting as it can only locate the origin ability developed during the training procedure. The LLM cannot learn the new in-context information with noisy labels. Instead, it only helps to locate the most similar concept seen during the pre-training stage. Despite the labels being noisy, ICL may still be able to locate the correct concept with the input text information. Empirical evidences Min et al. (2022) indicates that even random permute the model label can lead to a satisfying performance.
E.3 Abstraction Ability of LLMs
Despite the success of LLM based in the natural language, Webb et al. (2023); Mirchandani et al. (2023); Huang et al. (2023b); Chen et al. (2023) indicate the effectiveness on abstract symbol without knowing semantic meanings of any individual symbol. Webb et al. (2023) exhibits the emergence ability of LLM for abstract pattern induction while Mirchandani et al. (2023) suggest that LLM is a general pattern machine extrapolating sequences of numbers that represent states over time to complete simple motions. Huang et al. (2023b) achieves comparable performance using random Gaussian vectors instead of the original token embedding when context is sufficient. Chen et al. (2023) indicates such abstraction with randomizing embeddings can help LLM learn multiple languages.
E.4 Discussion On the Self-correction
The self-correction Pan et al. (2023); Kim et al. (2023); Gou et al. (2023); Welleck et al. (2022) is an advanced ICL technique iteratively revise the outputs of LLM utilizing feedbacks, aiming to mitigate undesired and inconsistent behaviors, e.g., lexically constrained generation and toxic reduction. Despite its effectiveness, the underlying mechanism remains an open question. The initial observations can be found as follows. Kadavath et al. (2022) illustrates positive evidence where LLM can accurately examine the correctness of their statements, serving as the necessary condition for self-correction. Nonetheless, Huang et al. (2023a) observes that self-correction cannot improve the performance since the added feedback may bias the model away from producing an optimal response to the initial prompt. Hong et al. (2023) provides more detailed evaluation setting and identifies that (1) LLMs perform much worse at identifying fallacies related to logical structure than those related to content. (2) LLMs cannot classify different types of fallacies. Despite the above phenomenons, there is still no understanding of the underlying mechanism of self-correction so far.
E.5 How The Data-generating Functions Are Different Than Arbitrary Functions
We first emphasize the importance of the data generation function. The strong generative capability is an essential ability for LLMs. Most successful applications and usage of the LLM revolve around the generative capability. Therefore, the data generation perspective is essential to understand the LLM.
The data-generating function is generally utilized to understand the data-generation capability of LLMs. It can be defined as ’the underlying hypothesis on textual data generation’. Technically, the data-generation function can be any function that can model the probability over a potential token given a sequence of tokens, after being trained with text data. The main difference between the data-generation function and arbitrary function is whether the function can be used to generate reasonable natural language sequences. Understanding the data-generation process is a core problem in natural language processing, particularly for natural language generation tasks.
More concretely, N-gram, HMM, and Recurrent Neural Networks are three straightforward data-generation functions but they cannot model long contexts, and the first two are non-parameterized data-generation functions. On the other hand, we can have a linguistic-driven data generation function, e.g. probabilistic context-free grammar Hahn and Goyal (2023), to introduce some priors of syntax. Since the complicated and hierarchical nature of human languages, LLMs are great in terms of incorporating contextual information through a powerful function approximation ability. Honestly speaking, we can claim that the impressive results of LLMs depend on the ability to approximate the unknown data-generation function underlying the pre-training corpora.
Notably, the statistical framework, which utilized the input-label mapping as the data generate function is a simplified setting. Such a simplified setting enables to conduct of more theoretical analysis. Therefore, we can qualitatively analyze the expressiveness, generalization, and internal mechanisms of the ICL. For instance, with the function abstraction, we can analyze the generalization within the same function class and between different function classes. However, how to take advantage of it in a real-world scenario remains unclear.
E.6 Whether Different Demonstrations Represent Different Data Generation Functions
Whether different demos represent different data generation functions depends on the hypothesis of the data generative function. It is possible for different demonstrations to share the same data generation function. On the contrary, it is also possible for different orders of the demonstrations to correspond to different data generation functions.
E.7 Whether there is the connection between skill learning/recognition and model under/overfitting?
The ICL procedure does not have any backward learning process, i.e. gradient descent, generally utilized in deep learning. Therefore, the ICL procedure is not explicitly related to the model under/overfitting without an explicit fitting procedure.
Both skill learning and skill recognition can achieve a certain generalization, without explicit under-fitting or over-fitting. The skill recognition is not directly memorization. Given the train data $(\mathbf{x},\mathbf{y})$ generated from the function $\mathbf{y}=\mathbf{k}\mathbf{x}$, the pre-training data can be within the input interval $\mathbf{x}\in[0,1]$, while the ICL test data can be within the input interval $\mathbf{x}\in[1,2]$. In such a case, the ICL can still achieve satisfying performance, indicating the generalization ability. It indicates the ICL with skill recognition can achieve generalization when test data are within the same function. A more comprehensive discussion when meeting out-of-distribution scenarios can be found in Appendix F.
The difference between skill learning and recognition is the different extent of the generalization. The skill recognition generalizes through seeking an existing function within the same function class but skill learning can come up with a new function within this function class.
E.8 The real-world correspondence of data generation functions
Our paper focuses on whether the ICL can learn a new data generation function in context. From a practice scenario, the new data generation function can be defined as the n-gram does not appear in the training stage. Such compositional generalization is a key concept in the NLP domain. For instance, such out-of-distribution can happen when LLMs read the news. The skill learning mechanism can learn the new n-gram and knowledge in context, while skill recognition tries to map the pre-training knowledge with the news.
Appendix F The Robustness of ICL On the Statistical Framework
We primarily analyze the skill-learning mechanism when (1) data generation functions during the pre-training and ICL inference stages are from the same function class, and (2) input features are sampled from the same distribution in Section 5. In this section, we provide a further discussion of how the skill-learning mechanism works when distribution shifts happen, indicating the robustness of the ICL. The robustness of the ICL is evaluated in different out-of-distribution scenarios, which can be roughly divided into the following categories: (1) Task shift, where the pre-training and in-context labels are generated from different function classes, is discussed in Appendix F.2. (2) Corvariate shift, where the pre-training and in-context inputs are sampled from different distributions, is discussed in Appendix F.3. (3) Query shift, where the in-context training inputs and the query sample input are sampled from different distributions, is discussed in Appendix F.4. Notably, all the above out-of-distribution scenarios are conducted on the statistical framework while it remains an unclear correspondence to the real-world LLM system pre-training on the massive corpus. More recently, Vladymyrov et al. (2024) focuses on the corrupted training data scenario with noises on different extend. Both empirical and theoretical results indicate the robustness of transformers in such scenario.
F.1 Preliminary
To formally describe different out-of-distribution scenarios, we first provide a rigorous description of the pre-training and prompt data from a distribution perspective. The pre-training data is defined as $(\mathbf{x}_{1},\mathbf{h}(\mathbf{x}_{1}),\cdots,\mathbf{x}_{N},\mathbf{h}(\mathbf{x}_{N}),\mathbf{x}_{\text{query}})$ where $\mathbf{x}_{i}\sim\mathcal{D}_{\mathbf{x}}^{\text{train}}$, $\mathbf{x}_{\text{query}}\sim\mathcal{D}_{\mathbf{x}}^{\text{train}}$ and $\mathbf{h}\sim\mathcal{D}_{\mathcal{H}}^{\text{train}}$. The test prompt is defined similarly but drawing from a different distribution where $\mathbf{x}_{i}\sim\mathcal{D}_{\mathbf{x}}^{\text{test}}$ and $\mathbf{x}_{\text{query}}\sim\mathcal{D}_{\mathbf{x}}^{\text{test}}$. We then describe different out-of-distribution scenarios and how the LLM behaves on them differently in the following sections.
F.2 Task Shift
Task shift Zhang et al. (2023a) is a concept shift which be formally defined as $\mathcal{D}^{\text{train}}_{\mathcal{H}}\neq\mathcal{D}^{\text{test}}_{\mathcal{H}}$. It describes that the pre-training and in-context labels are generated from different function groups. Existing literature demonstrates two different task shifts, i.e., noise shift Zhang et al. (2023a), and regression vector shift Raventos et al. (2023).
Noise shift Zhang et al. (2023a) corresponds to the scenario where the shift is induced by the random Gaussian noise. Typically, the pre-training data generation function is $\mathbf{y}=\left\langle\mathbf{w},\mathbf{x}\right\rangle$ where in-context data generation function is from noisy linear function $\mathbf{y}_{i}=\left\langle\mathbf{w},\mathbf{x}\right\rangle+\mathbf{\epsilon}$. Zhang et al. (2023a) observes satisfying performance under such shift, indicating the robustness under such Gaussian noise.
Regression vector shift Raventos et al. (2023) corresponds to the scenario where pre-training data generation functions are a limited group $\mathcal{F}_{\text{train}}$ of linear functions $\mathbf{f_{i}}:\mathbf{y}=\left\langle\mathbf{w}_{i},\mathbf{x}\right\rangle+\mathbf{b}_{i}$, where $\mathbf{f_{i}}\in\mathcal{F}_{\text{train}}$ The in-context data generation function is from all the possible linear functions covering the entire function space $\mathbf{f}_{i}\in\mathcal{F}_{\text{context}}$, where $\mathcal{F}_{\text{train}}\subseteq\mathcal{F}_{\text{context}}$. The task shift appears on the unseen data generation function during training. Raventos et al. (2023) observes that ICL exhibits the generalization gap with insufficient pre-training data. The emergence happens when the number of pre-training functions increases with satisfying out-of-distribution performance.
F.3 Covariate Shift
Covariate shift Zhang et al. (2023a) can be formally defined as $\mathcal{D}^{\text{train}}_{\mathbf{x}}\neq\mathcal{D}^{\text{test}}_{\mathbf{x}}$. It describes that the pre-training inputs and the in-context inputs are sampled from different distributions. Existing literature demonstrates different covariate shifts including low-dimensional subspace shift, skewed covariance shift, mean shift, and random covariate shift.
Low-dimensional subspace shift Garg et al. (2022) samples prompt input feature from random 10-dimensional subspace from the pre-training input feature. Garg et al. (2022) empirically observes the robustness over such covariate shift.
Skewed covariance shift Garg et al. (2022) samples in-context features from $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$ where $\mathbf{\Sigma}$ is a skewed covariance matrix with eigen-basis chosen uniformly at random and $i^{\text{th}}$ eigenvalue proportional to $1/i^{2}$. Empirically observations Garg et al. (2022) indicate the performance degradation when the input feature dimension is larger than 10.
Mean shift Ahuja and Lopez-Paz (2023) samples train and test inputs from $\mathcal{N}(\mathbf{\mu}_{\text{train}},\mathbf{\Sigma})$ and $\mathcal{N}(\mathbf{\mu}_{\text{test}},\mathbf{\Sigma})$ where $\mathcal{N}(\mathbf{\mu}_{\text{train}}\neq\mathcal{N}(\mathbf{\mu}_{\text{test}})$. Despite performance degradation to a certain extend, the transformer backbone shows better generalization than the MLP backbone with both empirical observations and theoretical evidence.
Random covariate shift Zhang et al. (2023a) corresponds to that pre-training training prompts and in-context prompts are sampled from distributions with different covariates. The ICL performance degradation Von Oswald et al. (2023); Zhang et al. (2023c) drops to 0 quickly with theoretical explanation Zhang et al. (2023c). The larger transformer with non-linearity serves as the solution to random covariate shift, while the reason underlying the emergent ability remains unclear.
F.4 Query Shift
Query shift Zhang et al. (2023a) is the covariate shift, which can be formally defined as $\mathcal{D}^{\text{test}}_{\text{query}}\neq\mathcal{D}^{\text{test}}_{\mathbf{x}}$. It describes the distribution shift within the in-context training samples and test samples are sampled from different distributions. Different from the task shift focusing on the distribution shift between pre-training data and prompt data, query shifts describe the distribution shift within the prompt data, where the training prompt data distribution is different from the prompt query distribution. Existing literature demonstrates two different query shifts as follows.
The orthants shift changes the positive or negative signs to each coordinate of in-context features, ensuring both prompt data and prompt query fall within the same orthant, distinct from the query input’s orthant. Garg et al. (2022) observes the robustness to this shift when differences between orthants are not large.
The orthogonal shift maps the the prompt query to the orthogonal space of prompt data, which is an extreme case of the formal one. Garg et al. (2022) shows empirical evidence where the prediction will be zero and the error will be significantly large. Zhang et al. (2023c) further theoretically underpins the underlying reason while no solution is found currently.